How do you get the negative of 4 to the 7th power?
Negative numbers to even powers will be positive, negative numbers to odd powers will be negative. The answer is -16,384.
36
What are the odd of roling any number with one dice?
When rolling a standard six-sided die, the odds of rolling any specific number (1 through 6) are 1 in 6, or approximately 16.67%. Since there are six possible outcomes, the probability of rolling any number is simply the ratio of favorable outcomes (1) to total outcomes (6). Thus, the odds of rolling any number are equal for each side of the die.
What is the only number not to have a reciprocal?
The only number to not have a reciprocal is zero. This is because it can't be divided by anything.
A pair of dice is rolled What is the problability fo getting a sum of 2?
Assuming you meant a standard pair of dice, only 1 and 1 add to 2, since there are no negatives, and numbers only stretch up to 6, and due to that, there is no 0. Therefore, 1/36 is the probability (probability of rolling a 1, 1/6, probability of rolling a 1, 1/6, 1/6(1/6) = 1/36).
How is it possible to not roll a 6 with 6 dice?
if you are talking about getting a roll that totals 6 - there is only one way to roll that - get all ones. as such there are 6^6 -1 = 46656 -1 = 46655 other ways to roll the dice. so the odds of not rolling a total of 6 with 6 dice is 46655/46656 = .~99.997857%.
If you are talking about not rolling a 6 on ANY of the dice, there are 5 ways to roll each die that will give you something other than 6 so the number of ways to not roll any 6's is 5^6 = 15625. That means the odds of not rolling any 6's on 6 dice is 15625/46656 = ~38.489798%
What is frequency approach to probability?
The frequency approach or experimental probability involves the estimation of the probability of an outcome as the proportion of "successful" outcomes in repeated trials.
I put "successful" in quotes because in epidemiology, a "successful" outcome is often a person catching a disease or infection and possibly dying as a result. I am not sure if anyone would consider that a success!
It is 0.3292, approx.
What the probability of rolling a sum of nine with a single toss of two fair six sided dice?
There are 36 different ways to roll 2 6-sided dice.
The ones that make 9 are:
3+6 , 4+5 , 5+4 , 6+3
4 out of 36 = 4/36 = 0,111111111 = 11,1 %
Out of the 36 possible and equally-probable outcomes (1st die's number and 2nd die's number), only 3 combinations give you a sum greater than 10:
* 1st die gives a 5 and 2nd die gives a 6
* 1st die gives a 6 and 2nd die gives a 5
* 1st die gives a 6 and 2nd die gives a 6
Therefore the probability that the sum is greater than 10 equals 3/36, or 1/12.
How many possible outcomes are there of choosing 1 number from the first 6 positive whole numbers?
Six.
When a coin is tossed why it appears head or tail?
Those are the names given to the two faces of a coin. From early days, coins had an image of important persons - kings, queens, emperors, on one face and that face was called heads. The reason for the other side being called "tails" is uncertain.
These are the only two outcomes of a coin being tossed. Although there is a non-zero probability of the coin falling and remaining on its edge, it is small enough to be ignored in practice.
What is the probability of rolling two fair value dice getting two different numbers?
Pr(Two different numbers) = 1 - Pr(Two same) = 1 - 1/6 = 5/6 = 83.3%
Are you more likely to have a boy or girl baby based on number of sperm realesed?
No, it makes no difference.
How To get valid results small samples are sufficient?
I will assume the sample is random. In general, the larger the sample, the smaller the percentage error will be (the difference between percentages in the sample, and the percentages in the universe from whence the sample is taken). The percentage error tends to go down as the square root of the size of the sample.
What is the ratio of aces to the total number of cards in a standard deck of cards?
1-14 stupid, seriously, how stupid are you?
What has the subject matter to do with it? It would help potential answerers to give an accurate reply if questions were precisely worded.
How many eyes are in a pack of 52 cards?
42, but maybe more.
In the traditional standard English deck design, there are 42 eyes
2 x 2 eyes on each of the 4 queens ==> 16 queen eyes
2 x 2 eyes on 3 kings, but 2 x 1 on the king of diamonds ==> 14 king eyes
2 x 2 eyes on the jacks of diamonds and clubs, but 2 x 1 on the jacks of hearts and spades ==> 12 jack eyes
But continental European decks have different designs for their royalty, which usually show all 4 eyes for each of the face cards, making a total of 48 eyes.